Verify that the given differential equation is an exact: 

(3x2+2y)dx+(x-4y)dy=0


1
Expert's answer
2022-01-04T11:26:00-0500

To verify if the given differential equation is exact,

Give the equation

"(3x^{2}+2y)dx+(x-4y)dy=0"

"where"

"M=3x^{2}+2y" and

"N=x-4y"

For exactness, we know that

"\frac{∂M}{∂y}=\frac{∂N}{∂x}"


Now, we shall find "\frac{∂M}{∂y}"

"\frac{∂M}{∂y}=2"

Also, we shall find "\frac{∂N}{∂x}"

"\frac{∂N}{∂x}=1"


Since, "\frac{∂M}{∂y} ≠ \frac{∂N}{∂x}"


Hence, the differential equation "(3x^{2}+2y)dx+(x-4y)dy=0"

Is not exact.